1) Calculate the energy required to produce one tonne (1000kg) of CaO from CaCO3.
2) How much natural gas is required to produce 1 tonne of CaO given that the heat of combustion of natural gas (heating value) is 22500 J/kg?
3) Calculate the amount of CO2 waste produced to make 1 tonne of CaO from CaCO3?
ΔHf data
CaCO3= -1206.9 kJ/mol
CaO = -635.09kJ/mol
CO2 = -393.51 kJ/mol
1
Expert's answer
2018-04-07T08:23:13-0400
1) CaCO3=CaO+CO2 ∆H0f = + 178.3 kJ/mol;
Q = 178.3*1000/56 = 3184 kJ.
2) CaCO3=CaO+CO2
∆H0C (natural gas) = 22500 J/kg
m (natural gas) = 3184 kJ / 22500 J = 0.14 kg.
3) n(CaO) = 17.86 mol = n (CO2)
m (CO2) = n * M(CO2) = 17.86 * 44 = 785.71 kg.